TThe personnel office at a large electronics firm regularly schedules job interviews and maintains records of the interviews. From the past records, they have found that the length of a first interview is normally distributed, with mean μ = 32 minutes and standard deviation σ = 7 minutes. (Round your answers to four decimal places.)

(a) What is the probability that a first interview will last 40 minutes or longer?


(b) Five first interviews are usually scheduled per day. What is the probability that the average length of time for the five interviews will be 40 minutes or longer?
Answer #1

a) P(Lasts 40 minutes or longer)

= P(z > (40 - 32)/7)

= P(z > 1.14)

= 0.1265

b) P(Mean 40 minutes or longer)

= P(Z > 40 – 32. 7/15

= P(z > 2.56)

= 0.0053